### v1

# Pairing in Fission: Studies of Static Approach and Collective Inertias

Abstract
Pairing plays a crucial role in the microscopic description of nuclear fission. Microscopic methods provide access to three quantities related to pairing, namely, the pairing gap ($\Delta$), the particle number fluctuations ($\Delta\hat{N}^{2}$), and the quenching factor (QF). The aim of this work is to analyse the impact of each of these quantities on the static description of the fission process.
spontaneous fission, potential energy surface, half-lives, pairing correlations, least-energy path, Gogny-type nuclear interaction, self-consistent methods
I Introduction
The nuclear fission process discovered more than 80 years ago Hahn and Straßmann (1939) still remains one of the most fascinating and not fully understood nuclear phenomenon Schmidt and Jurado (2018); Bender et al. (2020). Shortly after the first observation of fission in the laboratory, the process was given an explaination based on the liquid drop model assumption of the nucleus Meitner and Frisch (1939) that is still valid nowadays. The subsequent discovery one year later of spontaneous fission Flerov and Petrzhak (1940), indicated that the process itself is much more complex than initially thought. These early studies opened up the door to a myriad of fascinating experimental and theoretical studies that have been carried out and published since then.
The non-negligible role of pairing correlations in fission dynamics strongly affects the calculation of half-lives. Two main factors play a crucial role in this relationship. First, modification of the fission barrier height and nuclear deformation at the barrier. Second, reduction of the collective inertia during the increase of deformation from ground state to scission. These aspects have been extensively studied in the literature, employing a large variety of nuclear models.
The first research on the impact of pairing interaction on fission was conducted in the 1970s when Ledergerber and Pauli Ledergerber and Pauli (1973) determined fission paths using the liquid drop model with Strutinski shell correction and Bardeen-Cooper-Schrieffer (BCS) surface pairing in the space of deformation, including asymmetric shapes. The authors showed that the value of pairing gap $\Delta$ in the BCS model strongly affects the inertia parameter. In consequence, pairing interaction leads to the reduction of the penetration integral at large deformation without modifying the potential energy.
Moreto and Babinet  Moretto and Babinet (1974) applied a simple, realistic model to show that an increase in the pairing gap parameter leads to an increase in the penetrability of the fission barrier. The main message from this study is that there is a strong interplay between the potential energy, which increases as a square of the pairing gap, and the collective inertia, which decreases as an inverse of the pairing gap squared. As a consequence, the least action path leads through regions with a larger pairing gap than the least energy path.
The influence of pairing interaction on fission observables has been continued in a series of papers of the group from Maria Curie-Sklodowska University in Lublin Staszczak et al. (1985); Łojewski and Pomorski (1980); Staszczak et al. (1989); Łojewski and Staszczak (1999). The fission half-lives have been calculated in various macroscopic-microscopic approaches with BCS pairing. Proton ($\Delta_{p}$) and neutron ($\Delta_{n}$) pairing gap parameters were treated on the same footing as deformation parameters. The dynamical fission path was calculated by minimizing the action integral in the multidimensional space that included pairing degrees of freedom Staszczak et al. (1985); Łojewski and Staszczak (1999). It was shown that in 252Fm, the shortest half-lives are obtained for the fission path with larger $\Delta$ than in the static path that minimizes energy. Despite the increased size of the fission barrier by 1.4 MeV, the mass parameter along the dynamic path was reduced by half, leading to the reduction of half-lives by three orders of magnitude. A similar effect was observed along the fermium isotopic chain.
The impact of the quadrupole pairing on the spontaneous fission half-lives of Fm isotopes has been studied in Łojewski and Pomorski (1980) in the static approach. The authors observed the decrease of the second barrier height by 1 MeV for the quadrupole pairing forces, whereas the mass parameters may increase, especially on the first fission barrier. In some Fm isotopes, including quadrupole together with the monopole pairing leads to a reduction of half-lives by two orders of magnitude.
In Ref. Staszczak et al. (1989), the generator coordinate method, together with the Gaussian overlap approximation (GCM+GOA), has been applied to investigate the influence of pairing vibrations on the spontaneous fission half-lives calculations of even-even transplutonium isotopes. The authors found that the action integrals calculated along the static fission paths are around 13% larger than the integrals obtained from dynamical trajectories. Again, the reduction of spontaneous fission half-lives is the result of the decrease of mass parameters, which compensates the increase of fission barriers. The dynamic paring allowed to reduce the calculated half-lives by several orders of magnitude.
The findings mentioned above from macroscopic-microscopic models have also been observed in self-consistent studies. The authors of Ref. Karatzikos et al. (2010) presented the study of spontaneous fission barriers of heavy and super-heavy nuclei within the RHB theory and Relativistic Mean Field (RMF) plus BCS methods. The authors obtained an increase in the absolute value of the pairing energy and a strong decrease in barrier height with the increase of the pairing $\Delta$ along the fission path relative to its ground-state value.
In Ref. Sadhukhan et al. (2014); Giuliani et al. (2014); Zhao et al. (2016); Rodríguez-Guzmán et al. (2023), the impact of pairing on fission paths was studied using the particle-number dispersion $\Delta\hat{N}^{2}$ as collective degree of freedom. These studies, which involved Skyrme, Gogny, and relativistic energy density functionals, showed that the inclusion of $\Delta\hat{N}^{2}$ strongly modifies the least-action paths, producing an overall reduction of spontaneous fission lifetimes compared to static calculations and dynamical calculations where only mass multipoles are considered as collective degrees of freedom. Additionally, it has been demonstrated that in some nuclei pairing as a dynamic degree of freedom restores the axial symmetry at the first fission barrier Sadhukhan et al. (2014); Zhao et al. (2016). Time-dependent GCM+GOA calculation employing $\Delta\hat{N}^{2}$ as collective degree of freedom have shown that pairing correlations play an important role in shaping the fission fragment yields in induced fission Zhao et al. (2021). Also, $\Delta\hat{N}^{2}$ has been employed in beyond-mean field studies exploring the role of dynamic pairing correlations in fission Rodríguez-Guzmán et al. (2023).
An alternative approach to controlling the pairing interactions has been proposed in the Refs. Kouno et al. (2022, 2021) The method for modification of the strength of pairing interaction was achieved by changing the $G$ parameter in the RMF+BCS calculations. It was shown that increasing the pairing strength up to 20% reduces the fission barrier by around 2.5 MeV in Pu isotopes due to the higher single-particle level densities around the Fermi surface at the saddle point. This result is in contradiction to the research presented above when dynamic fission paths are characterized by stronger pairing and higher barriers at the same time. Recently, this method was intensively used in the analysis of the pairing interaction during the fission.
A similar relationship was found in Refs. Guan et al. (2021, 2023), where the authors employed a macroscopic-microscopic model with BCS pairing. It has also been found that the ground states and the scission configurations are less sensitive to the pairing strength than the saddle points. These results were confirmed in energy density functional studies employing the Skyrme forces in the fission of 240Pu Chen et al. (2022). The authors observed a decrease of the fission barriers up to 2.3 MeV in both symmetric and asymmetric fission channels when the pairing strength is increased by 10%. The deformation of a nucleus at the scission line can also be slightly modified by variation of pairing strength. A similar result has been presented in Ref.  Wang et al. (2023).
In microscopic Hartree-Fock-Bogoliubov (HFB) calculations, the impact of pairing correlations has also been studied by modifying the strength of the pairing field Giuliani and Robledo (2013); Rodríguez-Guzmán and Robledo (2014a, b); Guzman and Robledo (2017). These studies have shown that increasing the pairing strength by 5%-10% can reduce the spontaneous fission lifetimes by several orders of magnitudes, particularly for those nuclei with broad and high fission barriers, mostly because of the strong dumping of the collective inertias. Moreover, when pairing is stronger, the mass parameters show smoother behavior without sharp peaks along the fission path.
As the literature review presented here shows, in the last fifty years, the role of pairing in fission has been extensively studied. The non-negligible impact of pairing in fission barriers and half-lives is currently well known. However, as various theoretical methods were used to modify the strength of the pairing interaction, results obtained from different studies cannot always be directly compared. Moreover, some contradictory conclusions were drawn from these studies, namely differences in the modification of the size of the fission barrier with pairing. These ambiguities indicate the need to systematize the knowledge regarding the impact of pairing in static fission calculations. In this work, we closely examine the three self-consistent model parameters corresponding to pairing and thoroughly investigate their impact on the description of the fission process, employing a consistent theoretical framework based on the HFB calculations. The paper is organized as follows: Section II briefly presents the theoretical model employed in this work. Section III presents the results of our studies. We first address the impact of the pairing gap, the particle number fluctuations, and the quenching factor parameters on fission barriers. We then discuss the results of fission half-lives for various pairing strengths and different aspects of collective inertia parameters along fission paths. In Section IV, we summarize our main conclusions. In Appendix A, we discuss the problem of the relative sign of the occupancies in the Bogoliubov transformation.
II Theoretical framework
The starting point for our work is the computation of potential energy surface (PES). We use the microscopic, self-consistent HFB model with the finite-range Gogny-type nuclear interaction. We applied the widely used and well-established D1S parametrization Gogny (1973, 1975); Dechargé and Gogny (1980); Berger et al. (1989, 1991). The computations were done in a one-center axially symmetric, deformed harmonic oscillator basis with 16 shells (including as many as 23 one-dimensional harmonic oscillator shells along the symmetry axis $n_{z\,\textup{max}}=22$). The spontaneous fission half-lives are computed using the Wentzel-Kramers-Brillouin (WKB) approximation, assuming the axially symmetric quadrupole moment $Q_{20}$ as the leading coordinate. The estimation of the probability of tunneling through the one-dimensional fission barrier requires the computation of the action $S$ along the least energy path Brack et al. (1972):
[…]
being in and out the classical turning points defined by $V(Q_{20})=E_{0}$. The potential energy in the expression of the action $V(Q_{20})$ is the sum of the HFB energy and the rotational and zero point energy correction Egido and Robledo (2004); Ring and Schuck (1980). The $E_{0}$ parameter represents the quantal ground state energy obtained by solving the collective Schrödinger equation with the quadrupole moment as a degree of freedom. Although this parameter can be easily estimated by assuming a parabolic behavior of the PES around the ground state minimum and considering the collective inertia at that point, we employ a more pragmatic approach and take it equal to 0.5 MeV. $B(Q_{20})$ is the collective inertia calculated within the perturbative cranking approximation Baran et al. (2011):
$B(Q_{20})=\frac{M_{-3}(Q_{20})}{M_{-1}^{2}(Q_{20})}\,,$ (2)
where $M_{-n}$ are the so-called moments of order $n$ of the collective coordinate $Q_{20}$ (see Baran et al. (2011) for the definition). The formula for the inertia is based on the assumption that the collective operator is a one-body one, which is not true for the fluctuation on the number of particles operator $\Delta\hat{N}^{2}$. To overcome this problem, the traditional approach is to take the one-body part in the expression of $\Delta\hat{N}^{2}$ in the quasiparticle basis. Although this seems a reasonable assumption, it is desirable to test it against the results obtained from a truly one-body operator, like the pairing gap one discussed below. Within the WKB approximation, the spontaneous fission half-life (in seconds) is given by Schunck and Robledo (2016):
$t_{sf}=2.86\times 10^{-21}(1+\exp(2S)).$ (3)
As mentioned in the Introduction I, the goal of the present work is to discuss the impact of pairing correlations on the quantities entering the action integral (2). To do so, our analysis covers the three main methods that have been traditionally employed to characterize pairing correlations strengths in fission: HFB calculations with constraints on the pairing gap parameter ($\Delta$), on the particle number fluctuation operator ($\Delta\hat{N}^{2}$), and the renormalization of the pairing interaction.
The pairing gap operator can be defined as $\Delta=\frac{G}{2}(\hat{P}+\hat{P}^{+})$, being $\hat{P}^{+}$ the Cooper pair creation operator, and $G=10/A$ (MeV), a parameter related to the typical strength of the monopole pairing interaction. By employing $\Delta$ as constraining operator, it is thus possible to control the amount of pairing correlations in HFB wave functions $|\Phi\rangle$. The main advantage of employing $\Delta$ a collective degree of freedom is its one-body nature, which allows a consistent estimation of collective inertias through Eq. (2). However, in order to employ $\Delta$ as a collective degree of freedom, specific numerical strategies must be followed in order to properly evaluate the expectation value, as detailed in the Appendix A.
As one of the consequences of the existence of pairing correlations in a HFB calculation is the spontaneous symmetry breaking of the particle number symmetry, it is possible to employ one of the associated order parameters as a measure of pairing correlations. The most straightforward order parameter is associated with the particle number fluctuation $\Delta\hat{N}^{2}=(\hat{N}-\langle\hat{N}\rangle)^{2}$, being $\hat{N}$ the particles number operator. This quantity is zero in the absence of pairing correlations and becomes relatively large (of the order of a few units) for strongly correlated systems. As mentioned before, a drawback related to the fluctuation in particle number is its two-body-operator character. However, owing to the gradient method used in the solutions of the HFB equation, two-body constraint operators can be handled on the same foot as one-body ones.
As a last measure of pairing correlations, we consider the so-called quenching factor (QF), a parameter that renormalizes the strength of the pairing interaction of the Gogny force. The QF is introduced by multiplying the HFB pairing field $\Delta^{(\tau)}$ ($\tau=p,n$) by a constant factor in each iteration of the iterative solution of the HFB equation. This QF parameter is equal to 1.0 for not-modified HFB fields and takes typical values around 1.1 for enhanced pairing. It can be pictured as an analog of the pairing strength $G$ of other simple models, like BCS. In calculations employing the QF,
pairing correlations can be characterized by the so-called particle-particle energy $E_{pp}$ for protons and neutrons, which in HFB formalism is given by:
$E_{pair}^{(\tau)}=\frac{1}{2}\rm{Tr}(\Delta^{(\tau)}\kappa^{(\tau)}),$ (4)
where $\Delta^{(\tau)}$ is the pairing field and $\kappa^{(\tau)}$ stands for the pairing tensor obtained from the solution of the HFB equations. Although this quantity is not the pairing energy of the system (i.e., the energy difference between the paired solution and the corresponding Hartree-Fock energy), one usually refers to it as “pairing energy", and we will follow in subsequent sections this custom.
III RESULTS
Figure 1: (a) The PES (in MeV) of 262Rf isotope in the ($Q_{20},\Delta$) plane. The least-energy path is plotted with a white solid line. The potential energy (b), the pairing energy (c), and the collective inertias (d) along the least-energy path (black solid line) and for three constant values of the pairing gap.
Figure 2: Similar as Fig. 1 but when $Q_{20}$ and $\Delta\hat{N}^{2}$ are considered as collective degrees of freedom.
Figure 3: The PES in quadrupole moment and the pairing strength factor along the least-energy path. The energy is plotted in relation to the ground state energy of the nucleus for QF=1.
Figure 4: The HFB energy with the ZPE corrections (upper panel a,d), the perturbative mass parameters (middle panel b,e) as a function of $Q_{20}$ along the least-energy path and the corresponding pairing energy (c), (f) of protons (solid line) and neutrons (dashed line) for the 262Rf and 260Fm isotope.
Figure 5: The logarithms of spontaneous fission half-lives of 250Cm, 260Fm and 262Rf isotopes obtained within different values of the paring strength and compared with the experimental data taken from Kondev et al. (2021) depicted as horizontal lines.
Figure 6: (a) The least-energy path profiles of 238Pu (solid lines) and 236U (dashed lines) for different values of the quenching factor (QF) and (b) the corresponding collective inertias $B_{22}$ along the least-energy paths.
Figure 7: The logarithms of spontaneous fission half-lives of uranium (a) and plutonium (b) isotopes obtained within different values of the paring strength and compared with the experimental data (black circles).
Our first goal is to understand the impact of the pairing gap, particle number fluctuations, and the imposed pairing strength on the spontaneous fission static description. In order to clarify the picture and to minimize the influence of the secondary effects by involving the parity breaking, we decided to focus first on the case of symmetric fission. We have chosen for our analysis the 262Rf isotope for which spontaneous fission half-life, as well as the fission fragment mass yields, are measured and available for comparison.
III.1 The pairing gap
As has already been emphasized in the Introduction, the first studies within a simple, realistic model regarding the consideration of pairing in the description of fission showed that an increase in the pairing gap parameter leads to an increase in the penetrability of the fission barrier. This effect is strictly related to the fact that the collective inertias decrease as the inverse square of the pairing gap Moretto and Babinet (1974). Panel (a) of Fig. 1 depicts the potential energy surface (PES) in the ($Q_{20},\Delta$) plane of 262Rf. The solid white line represents the least-energy path in this space of the collective coordinates. To facilitate an understanding of the influence of $\Delta$ on the potential energy of the nucleus, we have also created a one-dimensional plot of the potential energy (panel b) and pairing energy (panel c) as a function of nuclear elongation. The black line corresponds to the path of least energy, while the other colors represent changes in energy as a function of $Q_{20}$ for a fixed value of the pairing gap greater than what was found in one-dimensional calculations with the constraint only on the quadrupole moment. It is evident that as $\Delta$ increases, the energy of the system increases as well. As expected, this is accompanied by an increase in the absolute value of pairing energy, as shown in panel (b), and a significant decrease in the values of collective inertias, as shown in panel (d). Apart from that, one should notice that the fission barrier becomes gradually broader for higher $\Delta$ values.
III.2 The particles number fluctuation
The particle number fluctuations $\Delta\hat{N}^{2}$, which is a quantity proportional to the pairing gap, has been successfully used for estimations of the spontaneous fission half-lives of actinides Sadhukhan et al. (2014); Giuliani et al. (2014); Zhao et al. (2016); Rodríguez-Guzmán et al. (2023). Most of the studies found that the coupling of $\Delta\hat{N}^{2}$ with shape degrees of freedom greatly improves the agreement between theoretical and experimental results. Now, we have decided to explore in detail how the particle number fluctuations as a collective degree of freedom impacts the fission barriers. The panel (a) of Fig. 2 presents the PES of 262Rf in the ($Q_{20},\Delta\hat{N}^{2}$) plane. One can notice that the topography of the presented surface looks very similar to Fig. 1. Also, the behavior of the analyzed quantities presented in panels (c) and (d) is similar to the one observed when the pairing gap was employed as a collective degree of freedom. This is a direct consequence of the almost linear connection between $\Delta$ and $\Delta\hat{N}^{2}$ variables. As $\Delta\hat{N}^{2}$ increases, the potential energy increases and the collective inertia decreases, both following a quadratic trend. As expected, in both cases, whether the pairing gap or the particle number fluctuations increases, it is accompanied by the broadening of the fission barrier As the energy and collective inertia profiles obtained with $\Delta$ and $\Delta\hat{N}^{2}$ show a very similar pattern, it is possible to conclude that these two quantities may be treated interchangeably.
III.3 The quenching factor
By considering different values for the quenching factor, one can scale the HFB proton and neutron pairing fields, which in practice leads to an increased or decreased pairing strength with respect to the value found in the process of fitting the parameters of the effective interaction. It has been shown that by increasing the quenching factor by 10-20%, one can obtain a better agreement with the data for the typical fission observable, like spontaneous fission half-lives of even-even Giuliani and Robledo (2013); Rodríguez-Guzmán and Robledo (2014a, b) and odd-mass Guzman and Robledo (2017) systems as well as fission fragment mass distributions Guan et al. (2023); Zhao et al. (2021). Fig. 3 depicts the least-energy path of 262Rf as a function of elongation for various values of the quenching factor. The presented PES is plotted in relation to the path obtained with $\textup{QF}=1$. Since the quenching factor is one of the three microscopic quantities related to pairing, one would expect that changes in the value of this parameter would lead to similar effects on energy and collective inertias as the previously analyzed $\Delta\hat{N}^{2}$ and $\Delta$. However, the results of the calculations conducted for the variable QF demonstrate that we are dealing with a different effect here, as we are changing the wave function by changing the underlying interaction. It should be noted that as the quenching factor increases, the potential energy of the nucleus decreases as the attractive pairing interaction becomes stronger. This is the opposite effect compared to what was observed earlier for increasing $\Delta\hat{N}^{2}$ and $\Delta$, which caused an increase in the potential energy. It is also worth noting that while increasing the particle number fluctuations, and the pairing gap results in higher and broader fission barrier, in the case of an increase in QF, the barrier becomes lower while its width does not change significantly. Furthermore, as shown in Fig. 3, further increasing the value of QF results in a monotonic decrease in the potential energy of the nucleus and a flattening of the fission barrier.
One can conclude that the quenching factor can be used as an artefact to renormalise locally the pairing strength in order to better reproduce experimental data. It can be also used for testing fission with enhanced pairing along the fission path. However, this method does not allow to dynamically adjust pairing strength for the search of the least action path.
III.4 The quenching factor and spontaneous fission half-lives evaluation
The usual approach to describe the spontaneous fission process is to follow the least-energy path. This is a reasonable approximation that produces qualitative and often quantitative results in good agreement with experimental data. By changing the quenching factor, one can expect a slightly different topography of the PES, as discussed in Sec. III.3. The upper panel of the Fig. 4 presents the fission barrier profiles corrected by the zero-point energy (ZPE) for 262Rf (left) and 260Fm (right). We notice a reduction of the fission barrier by a few MeV when the pairing strength is increased by 10%. The collective inertias $B_{22}$ are plotted in the middle panels of Fig. 4. Their values decrease together with increasing the quenching factor and become smoother. With the stronger pairing, the difference in the $B_{22}$ values becomes less apparent. The bottom panel of Fig. 4 presents the pairing energies of protons and neutrons. Obviously, the increase of the quenching factor leads to an increase of the absolute value of pairing energy. Fig. 5 presents the logarithms of spontaneous fission half-lives resulting from the static studies for various values of QF. As expected from the above analysis, the decrease of the fission barriers and the collective inertias leads to an increase of the probability of tunneling the fission barrier, leading to shorter half-lives as the pairing strength grows.
Let us now take a look at the actinides in order to have some insight into the pairing properties at the asymmetric fission path. For our studies, we have chosen the plutonium and uranium isotopic chains, where the asymmetric fission mode exists. At the panel (a) of Fig. 6, the least energy paths of two representative isotopes, namely 238Pu and 236U, are plotted for various values of QF. Here, the first hump of the barrier is reflection symmetric, whereas the second one is octupolly deformed. One can see that the increased QF lowers the first symmetric fission barrier by approximately 0.4 MeV per 0.05 change in the QF value. The second minimum and the second asymmetric fission barrier show a slightly smaller influence of pairing strength and the diminishing barrier height between the considered QF values does not exceed 0.2 MeV. This trend is inversed behind the second barrier, at around $Q_{20}$ = 80 b. Beyond this point, the barrier becomes higher for a higher value of the quenching factor. The collective inertias related to the discussed fission paths are presented in the lower panel of Fig. 6. Similarly, as for the previously discussed symmetric fission, $B_{22}$ is reduced and smoothened out for higher QF values. It is worth to mention that in these calculations, despite the second barrier in this region of actinides leads through the octupole shapes, the components of the inertia tensor associated with this degree of freedom have been neglected.
The calculated spontaneous fission half-lives are plotted in Fig 7. Looking at the results, one can immediately notice that the optimal value of the pairing strength for reproducing experimental half-lives seems to be between 105% and 110% for the D1S parametrization. The effect of the increased tunneling probability when the quenching factor increases, and consequently the reduction of the spontaneous fission half-lives, mostly comes from the action integral of the first hump of the fission barrier. This is because of the strong reduction of collective inertias with higher QF, accomplished by diminishing the potential energy. The influence of the asymmetric part of the barrier on the half-lives with increasing QF is less important. First, despite this part of the barrier being relatively broad, the collective inertia is much smaller than at the first barrier. Second, the reduction of the potential energy around the peak is compensated by an increase of the energy in the tail of the barrier. It means that the action integral is insensitive to the energy changes introduced by the QF in the asymmetric part of the least-energy path. This suggests that QF has a stronger impact when the parity is not broken. Thus, the overall trend of the shorter half-lives for higher QF in the actinides mainly comes from the response of the symmetric barrier on the pairing strength value.
III.5 Collective inertias with constraints on $\Delta$ and $\Delta\hat{N}^{2}$
Figure 8: The effective inertia $B(s)$ along the quadrupole moment $s=Q_{20}$ for several values of $\Delta$ (solid line) and the corresponding $\Delta\hat{N}^{2}$ (dashed line) for 262Rf - see the text for detailed explanation.
Further investigation of the role of microscopic pairing in fission should involve a dynamic approach requiring the calculation of the collective inertia for the pairing modes. Among the available quantities associated with the pairing correlations discussed in this paper, the quenching factor can not be considered a dynamical degree of freedom as it involves the change of the interaction itself, and it is, therefore hardly consistent with the traditional frameworks to define collective inertias. On the other hand, the $\Delta$ and $\Delta\hat{N}^{2}$ collective variables can be used for that purpose, and it is interesting to compare the inertias associated with both of them. As already mentioned in Section II, $\Delta\hat{N}^{2}$ is a two-body operator, which means that the standard cranking approximation cannot be applied straightforwardly. Following the traditional procedure, we decided to check the collective inertia associated with this degree of freedom when one uses in the formulas for the inertia the quasiparticle one body part of $\Delta\hat{N}^{2}$. In order to check the validity of this assumption, we compared the resulting $\Delta\hat{N}^{2}$ inertias with those associated to $\Delta$. The effective inertia is defined in terms of the collective inertia tensor $B_{ij}$ along the fission path $s=Q_{20}$:
$B_{\rm eff}(s)=\sum_{i,j}B_{ij}\frac{dq_{i}}{ds}\frac{dq_{j}}{ds},$ (5)
where $B_{ij}$ is the inverse of the collective mass tensor calculated within the perturbative cranking approximation Baran et al. (2011), and $(q_{i},q_{j})$ stand for the collective variable pair $(Q_{20},\Delta)$ and $(Q_{20},\Delta\hat{N}^{2})$. We have considered several fixed values of $\Delta$ and related to their values of the $\Delta\hat{N}^{2}$ for which the collective inertias along the quadrupole moment have been computed. The results for 262Rf are shown in Fig. 8. The solid lines represent the inertias associated with $\Delta$ while the dashed lines correspond to the ones obtained when $\Delta\hat{N}^{2}$ is treated as collective variable. As one can see, the effective inertias associated with $\Delta\hat{N}^{2}$ agree very well with the $\Delta$-dependent ones. It is worth emphasizing that the behavior of both inertias becomes more consistent as the value of $\Delta$ and $\Delta\hat{N}^{2}$ increases. This is particularly important in dynamic calculations, as one should expect that the path of the least-action path will primarily pass through higher values of these constraints. These results suggest that, in first approximation, one can apply the one-body expression for the inertias in the perturbative cranking approximation employing the quasiparticle one body part of $\Delta\hat{N}^{2}$ as a dynamical variable. Developments in this direction are currently in progress.
IV CONCLUSIONS
In this work, we analyzed the different methods traditionally employed in studies exploring the impact of pairing correlations in the static description of spontaneous fission. We compared the potential energy surfaces, fission barrier, collective inertias, and pairing energies obtained when $\Delta$ and $\Delta\hat{N}^{2}$ are employed as collective degrees of freedom. The results have been confronted with those obtained when a quenching factor QF is applied to the HFB pairing field.
The obtained results of the microscopic studies of the impact of pairing on fission properties allow us to draw the following conclusions.
The strength of pairing interaction along the fission path from ground state to scission can be controlled in various ways. One may increase the absolute value of the pairing energy by raising the pairing gap ($\Delta$), the particle number fluctuations ($\Delta\hat{N}^{2}$), or the pairing strength.
The increase of $\Delta$ and $\Delta\hat{N}^{2}$ leads to an increase of the potential energy, accompanied by a decrease of the collective inertias. In consequence, higher and broader fission barriers are obtained. Despite this, the smaller inertia reduces half-lives by a few orders of magnitude.
Conversely, the increase of the pairing strength by scaling the quenching factor leads to a decrease of the potential energy. Lowering the fission barrier and the collective inertias simultaneously leads to a monotonic reduction of the fission half-lives for stronger pairing interactions. The effect of lowering the barrier with an increase of quenching factor is less significant in the region of the asymmetric fission channel. Thus, enhancing the pairing strength has a different impact on fission properties compared to the two previous techniques. This is because the nuclear interactions are modified. Changes in $\Delta$ and $\Delta\hat{N}^{2}$ touches only the single-particle level properties around the Fermi level, but the nuclear force remains unchanged.
The effective inertias associated with the $\Delta\hat{N}^{2}$ computed within the “one-body assumption" are in a very good agreement with the ones obtained within a standard formula applicable for one-body operators, giving a solid ground to pursue the comparative dynamical studies.
Static calculations of fission half-lives, taking into account various values of the quenching factor, do not exhaust the problem of describing spontaneous fission in terms of pairing forces. A more suitable treatment involves considering fission dynamics and should lead to a further reduction of computed spontaneous fission half-live. Work along these lines is already in progress. A similar approach should be applied to study fission-fragments mass yields. This would require, however, the inclusion of the octupole moment as an additional degree of freedom, which increases the numerical complexity.
Acknowledgements.
This research is part of the project No. 2021/43/P/ST2/03036 co-funded by the National Science Centre and the European Union’s Horizon 2020 research and innovation program under the Marie Skłodowska-Curie grant agreement no. 945339. The work of LMR and SAG is supported by Spanish Agencia Estatal de Investigacion (AEI) of the Ministry of Science and Innovation (MCIN) under Grant No. PID2021-127890NB-I00. SG acknowledges support funded by MCIN/AEI/10.13039/501100011033 and the “European Union NextGenerationEU/PRTR” under grant agreement No. RYC2021-031880-I.
Appendix A On the sign of occupation factor
The constraint in the gap parameter $\Delta=G\sum_{k}u_{k}v_{k}$ used in this paper presents a drawback associated with an overlooked property of the HFB and BCS theories. To discuss the problem, the BCS method will be used as the HFB one can be reduced to the former by working on the canonical basis of the density matrix. In the standard BCS theory, the canonical transformation to quasiparticles is defined as:
$\alpha_{k}^{+}$ $=u_{k}a^{+}_{k}-v_{k}a^{+}_{\bar{k}}\,,$ (6)
$\alpha_{\bar{k}}^{+}$ $=u_{k}a^{+}_{\bar{k}}+v_{k}a^{+}_{k}\,,$ (7)
where the variational parameters are assumed to be real and preserve time reversal invariance. The condition for BCS to be a canonical transformation requires that $u_{k}^{2}+v_{k}^{2}=1$. As the relative sign of the $u_{k}$ and $v_{k}$ coefficients in the BCS theory is not determined, traditionally, one chooses positive values for both sets of quantities. The most likely reason why the relative sign between $u$ and $v$ has never been considered is because pairing is often treated along with constant pairing interaction, for which all the relevant matrix elements are state-independent and attractive. In this specific case, the relative sign is irrelevant as the energy depends Ring and Schuck (1980) upon powers of $v^{2}$ (density matrix) or in the parameter $\Delta$ defined before. As the $\Delta$ contribution to the energy is quadratic and attractive, the minimization principle favors the largest possible values of $\Delta$ and, therefore, coherence in the relative sign of $u$ and $v$ (as incoherent signs will average to zero in $\Delta$). The same argument also applies to state-dependent pairing, as the matrix elements are predominantly attractive. However, if pairing matrix elements depend on the quantum numbers of the involved orbitals (say $k$ or $l$), it might happen that some products $u_{k}v_{k}$ give a larger contribution to the energy than others, $u_{l}v_{l}$. If we combine this possibility with a constrain in $\Delta$, it might happen that in order to increase the absolute value of $u_{k}v_{k}$ to gain energy and, at the same time, preserve the value of $\Delta$, some of the remaining terms $u_{l}v_{l}$ become negative. In our calculations preserving axial symmetry, $K$ (the projection of angular momentum along the $z$ axis) is a good quantum number, and all physical quantities, including the density and pairing tensor, are block diagonal in $K$. As a consequence, $\Delta$ receives independent contributions from each $K$ block, and we have witnessed in our calculations situations where all the $v_{k}$ of a given $K$ block change sign and give a negative contribution to $\Delta$. This configuration with the opposite sign of $v_{k}$ has lower energy than the one with all the $v_{k}$ having the same sign but obviously differ in the value of $\Delta$, and therefore they cannot be compared. In order to define $\Delta$ as a well-behaved collective variable, the sign of the $v_{k}$ has to be kept consistent for all the configurations. This difficulty does not show up obviously in the calculations where $\Delta\hat{N}^{2}$ is constrained as the particle number fluctuation depends on the square of both $u_{k}$ and $v_{k}$ and the sign degree of freedom is not relevant to determine the value of the constraint. Therefore, in order to circumvent the problem with the constraint in $\Delta$ we have followed a simple receipt: Do the calculation with a constraint on $\Delta\hat{N}^{2}$, change the relative sign between $u_{k}$ and $v_{k}$ to be positive and compute the almost linear relation $\Delta(\Delta\hat{N}^{2})$ to establish the correspondence between $|\Phi(\Delta\hat{N}^{2})\rangle$ and $|\Phi(\Delta)\rangle$. Such a linear dependence of $\Delta(\Delta\hat{N}^{2})$ for 262Rf in the relevant range of $\Delta$ is presented in Fig. 9.
Figure 9: The relation between $\Delta$ and $\langle\Delta N^{2}\rangle$ for 262Rf.
Please remember that the introduction of the constraint on $\Delta$ is to have a constraint on a one-body operator where the traditional collective inertia formulas are justified, as opposed to the case of the two-body constraint in $\Delta\hat{N}^{2}$.
We point out that BCS wave function differing in the sign of the $v_{k}$ do no longer have overlap one. The overlap between such wave functions is given by the product extended to all quantum numbers in the basis of the factors $u_{k}u^{\prime}_{k}+v_{k}v^{\prime}_{k}$. For those quantum numbers where the occupancy $v^{\prime}_{k}$ has the same sign as $v_{k}$, the factor is $u_{k}u_{k}+v_{k}v_{k}=1$. However, for those with opposite sign, the factor becomes $u_{k}u_{k}-v_{k}v_{k}\leq 1$, and therefore they contribute to decrease the overlap.